OpenStudy (anonymous):

A circle has its center at the center of a square with 26-inch sides. Using 3.14 for pi, find the area of the square not covered by the circle, then write a ratio of the circle’s circumference to the perimeter of the square.

OpenStudy (anonymous):

area of square = s^2 = (26)^2 area of circle = pi*r^2 = 3.14*(13)^2 area of square not covered by the circle = 26^2-3.14(13)^2 perimeter of square = 4s = 4(26) = 104 circumference of circle = pi*d = 3.14*26 circumference:perimeter ratio => 3.14:4